OpenStudy (anonymous):

I don't get trigonometric limits.. 1. Use the squeeze theorem to evaluate the limit as x approaches 0 of x^2sin(1/x) 2. limit as h approaches 0 of [sin(5h)]/3h 3. limit as x approaches 0 of (x^2)/sin^2(x) 4. limit as t approaches 0 of [cos(t)-cos^2(t)]/t 5. Use the squeeze theorem to evaluate the limit as x approaches 1 of (x-1)sin(pi/x-1) Thank you!

OpenStudy (anonymous):

one Question at one post 1) the greatest value of sin(something) is 1 and the smallest value is -1. squeeze sin(something) between -1 and 1

OpenStudy (anonymous):

\[ -1 \leq \sin(1/x) \leq +1 \\ - x^2\leq x^2\sin(1/x) \leq x^2 \] then put limits on both sides \[ \lim_{x->0} - x^2\leq \lim_{x->0}x^2\sin(1/x) \leq \lim_{x->0}x^2 \\ 0\leq \lim_{x->0}x^2\sin(1/x) \leq 0\]